# Exploring the Frontiers of Authorship Attribution in The Era of LLMs

### v1

@@ Abstract.
Accurate attribution of authorship is crucial for maintaining the integrity of digital content, improving forensic investigations, and mitigating the risks of misinformation and plagiarism. Addressing the imperative need for proper authorship attribution is essential to uphold the credibility and accountability of authentic authorship. The rapid advancements of Large Language Models (LLMs) have blurred the lines between human and machine authorship, posing significant challenges for traditional methods. We presents a comprehensive literature review that examines the latest research on authorship attribution in the era of LLMs. This survey systematically explores the landscape of this field by categorizing four representative problems: (1) Human-written Text Attribution; (2) LLM-generated Text Detection; (3) LLM-generated Text Attribution; and (4) Human-LLM Co-authored Text Attribution. We also discuss the challenges related to ensuring the generalization and explainability of authorship attribution methods. Generalization requires the ability to generalize across various domains, while explainability emphasizes providing transparent and understandable insights into the decisions made by these models. By evaluating the strengths and limitations of existing methods and benchmarks, we identify key open problems and future research directions in this field. This literature review serves a roadmap for researchers and practitioners interested in understanding the state of the art in this rapidly evolving field. Additional resources and a curated list of papers are available and regularly updated at https://llm-authorship.github.io/.
@@ 1. Introduction
Authorship Attribution (AA) is the process of determining the author of a particular piece of writing, which has significant real-world applications across various domains. In forensic investigations, authorship attribution plays a crucial role in solving murder cases disguised as suicides (Chaski, 2005; Grant, 2020), tracking terrorist threats (Winter, 2019; Cafiero and Camps, 2023), and aiding general criminal investigations (Koppel et al., 2008; Argamon, 2018; Belvisi et al., 2020). In the digital realm, authorship attribution helps safeguard the integrity of content by preventing deceptive social media activities (Hazell, 2023), detecting account compromises (Barbon et al., 2017), and linking user profiles across various social networks (Shu et al., 2017; Sinnott and Wang, 2021). Additionally, authorship attribution techniques are instrumental in combating misinformation (Chen et al., 2022b; Chen and Shu, 2024b, a; Shu et al., 2020; Stiff and Johansson, 2022; Hanley and Durumeric, 2024), protecting intellectual property rights (Meyer zu Eissen et al., 2007; Stamatatos and Koppel, 2011), and identifying fraudulent practices such as fake reviews (Ott et al., 2011; Afroz et al., 2012).
@@
The rapid advancements in LLMs have significantly improved text generation, producing outputs that rival human writing in fluency and coherence. This progress underscores the imperative need to distinguish between human-written text, LLM-generated text, or a combination of both. As illustrated in Figure 1, authorship attribution can be systematically categorized into four problems: attributing unknown texts to human authors, detecting LLM-generated texts, identifying the specific LLM or human responsible for a text, and classifying texts as human, machine, or human-LLM co-authored. Each task presents unique challenges that necessitate corresponding solutions. Researchers continually adapt and refine attribution methods, transitioning from human-authored texts to LLM-generated content, and navigating the complex interweaving in human-LLM co-authored works. As detection methods advance, adversarial attacks also evolve to bypass these measures, creating a continuous cycle of challenge and response in the quest to distinguish and disguise authorship (Dugan et al., 2024). Addressing these challenges will pave the way for more robust and reliable authorship attribution techniques.
Authorship attribution for both human and LLM-generated texts can be framed as either binary or multi-class classification. The LLM-generated text detection task simplifies the attribution problem by classifying each text as either originating from humans or LLMs (Jawahar et al., 2020a; Mitchell et al., 2023; Pu et al., 2023a; Sadasivan et al., 2023). The majority of previous research on the automatic detection of machine-generated text has focused on binary classification (Jawahar et al., 2020b; Mitchell et al., 2023). In the more challenging multi-class task, the goal is not only to differentiate between human and LLM-generated text but also to classify the text according to its specific source of generative models (Uchendu et al., 2021; Li et al., 2023b). Differences in LLM architectures, training methods, and generation techniques can influence the style of the generated texts (Munir et al., 2021). In the more complex human-LLM co-authoring problem, the goal is to distinguish texts authored by humans, LLMs, or combinations of both. Such nuanced detection provides deeper insights into the provenance of the text and is crucial for applications requiring detailed source attribution. Neural network-based detectors, generally outperform metric-based methods in both human authorship attribution and LLM-generated text detection problems (He et al., 2023; Zhang et al., 2024). However, these neural network approaches often offer less explainability compared to their metric-based counterparts.
@@
  • We provide a timely overview to discuss the challenges and opportunities presented by LLMs in the field of authorship attribution. By systematically categorizing authorship attribution into four problems and balancing problem complexity with practicality, we reveal insights into the evolving filed of authorship attribution in the era of LLMs.
@@ 2.2. Methodologies
Machine learning approaches integrate stylometric features with classifiers such as logistic regression (Aborisade and Anwar, 2018; Madigan et al., 2005a), Bayesian multinomial regression (Argamon et al., 2009), and support vector machines (SVM) (Bacciu et al., 2019). Before the widespread adoption of transformer-based models, multi-headed Recurrent Neural Networks (RNNs) (Bagnall, 2015), and Long Short-Term Memory (LSTMs) were utilized at both sentence and article levels (Qian et al., 2017). Convolutional Neural Networks (CNNs) were applied at various levels, including characters, words, and N-grams (Ruder et al., 2016; Shrestha et al., 2017a, b). Moreover, syntax-augmented CNN models (Zhang et al., 2018), Convolutional Siamese Networks (Saedi and Dras, 2021), and attention-based Siamese Networks (Boenninghoff et al., 2019a) were explored. Additionally, combinations of convolutions and transformers have been employed to learn embeddings for comparison tasks (Andrews and Bishop, 2019).
@@
Barlas and Stamatatos (2020a) found that BERT performed well with large vocabularies, outperforming multi-headed RNNs. Fabien et al. (2020) fine-tuned a BERT model for authorship attribution, showing that including additional stylometric and hybrid features in an ensemble model can improve performance. Rivera-Soto et al. (2021) concluded that topic diversity and dataset size are crucial for effective cross-domain transfer. Adaptation through style transfer has not resolved cross-domain issues (Boenninghoff et al., 2019b; Wegmann et al., 2022). Techniques like slanted triangular learning rates and gradual unfreezing can be used to avoid catastrophic forgetting during fine-tuning (Howard and Ruder, 2018).
@@
The incorporation of LLMs in authorship attribution addresses several limitations of traditional methods. Unlike BERT-based models, which require computationally expensive fine-tuning and large amounts of domain-specific data for optimal performance, LLMs can generalize across various domains without fine-tuning, thereby mitigating issues related to domain specificity (Barlas and Stamatatos, 2020a). LLMs are also effective with shorter texts, reducing the necessity for long inputs to derive meaningful representations (Eder, 2015). Another key advantage of LLM-based approaches is their ability to provide natural language explanations for their predictions, enhancing transparency compared to hidden text embeddings (Huang et al., 2024). This versatility marks a step forward in overcoming challenges related to data, domain specificity, text length requirements, and explainability faced by earlier methods.
@@ 3. LLM-generated Text Detection
LLMs excel at generating fluent and coherent text, raising concerns about the authenticity and originality of authorship. Detecting LLM-generated text is crucial for several applications, including combating misinformation and disinformation on social media (Gambini et al., 2022; Stiff and Johansson, 2022; Chen and Shu, 2024b), identifying spam (Jindal and Liu, 2008), preventing phishing attacks (Hazell, 2023), identifying fake reviews (Salminen et al., 2022), and detecting machine-generated scientific papers (Rodriguez et al., 2022a; Liu et al., 2023a). As a result, the detection of LLM-generated text has garnered significant attention (Kumarage and Liu, 2023; Tang et al., 2023a; Wu et al., 2023b; Yang et al., 2023b).
@@ 3.1. Problem Definition
LLM-generated texts are included within the scope of machine-generated texts [^1] Machine-generated texts is also referred to as machine-authored, AI-generated, neural-generated, deepfake text, neural text, or synthetic text.. Machine-generated texts encompass any text produced by automated systems, including simpler language models or rule-based systems (Uchendu et al., 2021). This paper focuses specifically on LLM-generated texts. The task of detecting LLM-generated text involves identifying and distinguishing text created by LLMs from that written by humans. Typically, this task is approached as a binary classification problem (Zellers et al., 2019; Solaiman et al., 2019; Jawahar et al., 2020b; Fagni et al., 2021; Mitchell et al., 2023).
@@ 3.2. Methodologies
Evaluating the quality of machine-generated excerpts has traditionally relied on human judgment, which is considered the gold standard for open-domain generation systems (van der Lee et al., 2019; Gehrmann et al., 2019a). However, distinguishing between LLM-generated and human-written texts poses significant challenges for humans (Dugan et al., 2023). For example, untrained human reviewers often fail to distinguish GPT-3-generated text from human-written text, identifying it correctly only at a rate consistent with random chance (Clark et al., 2021). Liu et al. (2023a) found that even experienced faculty members and researchers could only achieve about a 50% success rate in identifying GPT-generated academic writings. In contrast, detection algorithms frequently outperform human reviewers in this task (Ippolito et al., 2020).
@@
LLM-generated text detectors can be categorized into metric-based and model-based methods (He et al., 2023), which are further divided into feature-based, neural network-based, zero-shot-based, and watermark-based methods. These detectors are also classified as either white-box or black-box, depending on their access to LLM weights (Tang et al., 2023b; Yang et al., 2023b). Watermarking-based methods typically fall under white-box detection, while proprietary models are restricted to black-box methods.
@@
Zero-shot detection methods are generally statistics-based, enabling the identification of LLM-generated text without additional training (Su et al., 2023b). Various statistical measures have been employed for detection, including entropy (Lavergne et al., 2008; Gehrmann et al., 2019a), perplexity (Beresneva, 2016; Hans et al., 2024), average log-probability score (Solaiman et al., 2019), fluency (Holtzman et al., 2020), and Zipf’s word frequency law (Zipf, 2016; Piantadosi, 2014) [^2] The frequency of a word decreases as its rank in a frequency-ordered list increases.. Additional methods leverage n-grams (Yang et al., 2023a), Uniform Information Density (UID) (Venkatraman et al., 2023), log rank information (Su et al., 2023c), and various linguistic features such as part-of-speech determiners, conjunctions, auxiliary relations, vocabulary, and emotional tone (Joulin et al., 2017; Tang et al., 2023b; Gehrmann et al., 2019b).
Various zero-shot detection methods provide distinct strategies to enhance both detection accuracy and efficiency. Gallé et al. (2021) introduced an unsupervised method that identifies over-appearance of repeated higher-order n-grams, distinguishing them from human-generated text. DetectGPT (Mitchell et al., 2023) relies on the observation that LLM-generated passages often fall into regions of negative curvature in log probability. (Bao et al., 2023) enhanced this approach and proposed Fast-DetectGPT, which increases efficiency by using conditional probability curvature. Additionally, some methods leverage LLMs themselves for text classification (Zellers et al., 2019; Solaiman et al., 2019). Different decoding methods are often applied to generate more diverse and less repetitive text, though these can also lead to hallucinations and less verifiable content (Shakeel and Jain, 2021; Guo et al., 2023). Fact-checking methods can mitigate these issues (Zhong et al., 2020; Schuster et al., 2020). Additionally, Krishna et al. (2024a) developed a detector that uses information retrieval to store LLM outputs in a database and search for semantically similar content to identify LLM-generated text, though this method raises privacy concerns regarding the storage of user conversations.
@@
Further advancements have improved the robustness, efficiency, and stealthiness of watermarking methods (Hou et al., 2023; Wu et al., 2023a; Zhao et al., 2023). However, there is a trade-off between watermark effectiveness and text quality, as more reliable watermarks require more extensive text modifications (Sadasivan et al., 2023). Additionally, watermarking presents challenges for proprietary LLMs and third-party applications due to the necessity of accessing the language model’s logits (Kirchenbauer et al., 2023). Watermark-based detection methodologies are also vulnerable to paraphrasing attacks (Sadasivan et al., 2023; Krishna et al., 2024a).
@@ 3.3. Open Challenges
Detecting LLM-generated texts is challenging due to their high quality and ability to mimic human writing styles. LLMs can emulate human writing so closely that they pose significant challenges to traditional stylometric techniques. They can incorporate complex narrative structures and varied vocabularies, making it difficult to distinguish between human and LLM-generated texts. The rapid evolution of LLMs further complicates detection, as newer versions exhibit different stylistic traits, making detection models quickly obsolete (Chakraborty et al., 2023b). LLM-generated text detectors often struggle to generalize to unseen domains encountered during training (Pu et al., 2023b; Rodriguez et al., 2022b; Li et al., 2023a) and tend to perform better on LLMs they were specifically trained on (Pu et al., 2023b; Chakraborty et al., 2023b; Li et al., 2023a).
Existing detectors also lack robustness to various factors, such as alternative decoding strategies (Ippolito et al., 2020), input sequence length (Solaiman et al., 2019), different prompts (Kumarage et al., 2023; Lu et al., 2023), repetition penalties (Fishchuk and Braun, 2023), and human edits (Zhang et al., 2024). Additionally, detectors are vulnerable to adversarial attacks, including homoglyph attacks (Gagiano et al., 2021; Macko et al., 2024), whitespace insertion (Cai and Cui, 2023), syntactic perturbations (Bhat and Parthasarathy, 2020a), synonym replacement (Kulkarni et al., 2023), and paraphrasing (Krishna et al., 2024b; Shi et al., 2024; Becker et al., 2023).

### v2

# Authorship Attribution in the Era of LLMs: Problems, Methodologies, and Challenges

@@ Abstract.
Accurate attribution of authorship is crucial for maintaining the integrity of digital content, improving forensic investigations, and mitigating the risks of misinformation and plagiarism. Addressing the imperative need for proper authorship attribution is essential to uphold the credibility and accountability of authentic authorship. The rapid advancements of Large Language Models (LLMs) have blurred the lines between human and machine authorship, posing significant challenges for traditional methods. We present a comprehensive literature review that examines the latest research on authorship attribution in the era of LLMs. This survey systematically explores the landscape of this field by categorizing four representative problems: (1) Human-written Text Attribution; (2) LLM-generated Text Detection; (3) LLM-generated Text Attribution; and (4) Human-LLM Co-authored Text Attribution. We also discuss the challenges related to ensuring the generalization and explainability of authorship attribution methods. Generalization requires the ability to generalize across various domains, while explainability emphasizes providing transparent and understandable insights into the decisions made by these models. By evaluating the strengths and limitations of existing methods and benchmarks, we identify key open problems and future research directions in this field. This literature review serves a roadmap for researchers and practitioners interested in understanding the state of the art in this rapidly evolving field. Additional resources and a curated list of papers are available and regularly updated at https://llm-authorship.github.io/.
@@ 1. Introduction
Authorship Attribution (AA) is the process of determining the author of a particular piece of writing, which has significant real-world applications across various domains. In forensic investigations, authorship attribution plays a crucial role in solving murder cases disguised as suicides (Chaski, 2005; Grant, 2020), tracking terrorist threats (Winter, 2019; Cafiero and Camps, 2023), and aiding general criminal investigations (Koppel et al., 2008; Argamon, 2018; Belvisi et al., 2020). In the digital realm, authorship attribution helps safeguard the integrity of content by preventing deceptive social media activities (Hazell, 2023), detecting account compromises (Barbon et al., 2017), and linking user profiles across various social networks (Shu et al., 2017; Sinnott and Wang, 2021). Additionally, authorship attribution techniques are instrumental in combating misinformation (Chen and Shu, 2024b, a; Shu et al., 2020; Stiff and Johansson, 2022; Hanley and Durumeric, 2024), protecting intellectual property rights (Meyer zu Eissen et al., 2007; Stamatatos and Koppel, 2011), and identifying fraudulent practices such as fake reviews (Ott et al., 2011; Afroz et al., 2012).
@@
The rapid advancements in LLMs have significantly improved text generation, producing outputs that rival human writing in fluency and coherence. This progress underscores the imperative need to distinguish between human-written text, LLM-generated text, or a combination of both. As illustrated in Figure 1, authorship attribution can be systematically categorized into four representative problems: attributing unknown texts to human authors, detecting LLM-generated texts, identifying the specific LLM or human responsible for a text, and classifying texts as human, machine, or human-LLM co-authored. Each task presents unique challenges that necessitate corresponding solutions. Researchers continually adapt and refine attribution methods, transitioning from human-authored texts to LLM-generated content, and navigating the complex interweaving in human-LLM co-authored works. As detection methods advance, adversarial attacks also evolve to bypass these measures, creating a continuous cycle of challenge and response in the quest to distinguish and disguise authorship (Dugan et al., 2024). Addressing these challenges will pave the way for more robust and reliable authorship attribution techniques.
Authorship attribution for both human and LLM-generated texts can be framed as either binary or multi-class classification. The LLM-generated text detection task simplifies the attribution problem by classifying each text as either originating from humans or LLMs (Jawahar et al., 2020; Mitchell et al., 2023; Pu et al., 2023a; Sadasivan et al., 2023). The majority of previous research on the automatic detection of machine-generated text has focused on binary classification (Jawahar et al., 2020; Mitchell et al., 2023). In the more challenging multi-class task, the goal is not only to differentiate between human and LLM-generated text but also to classify the text according to its specific source of generative models (Uchendu et al., 2021; Li et al., 2023b). Differences in LLM architectures, training methods, and generation techniques can influence the style of the generated texts (Munir et al., 2021). In the more complex human-LLM co-authoring problem, the goal is to distinguish texts authored by humans, LLMs, or combinations of both. Such nuanced detection provides deeper insights into the provenance of the text and is crucial for applications requiring detailed source attribution. Neural network-based detectors, generally outperform metric-based methods in both human authorship attribution and LLM-generated text detection problems (He et al., 2023; Zhang et al., 2024). However, these neural network approaches often offer less explainability compared to their metric-based counterparts.
@@
  • We provide a timely overview to discuss the challenges and opportunities presented by LLMs in the field of authorship attribution. By systematically categorizing authorship attribution into four representative problems and balancing problem complexity with practicality, we reveal insights into the evolving filed of authorship attribution in the era of LLMs.
@@ 2.2. Methodologies
Machine learning approaches integrate stylometric features with classifiers such as logistic regression (Aborisade and Anwar, 2018; Madigan et al., 2005a), Bayesian multinomial regression (Argamon et al., 2009), and support vector machines (SVM) (Bacciu et al., 2019). Before the widespread adoption of transformer-based models, multi-headed Recurrent Neural Networks (RNNs) (Bagnall, 2015), and Long Short-Term Memory (LSTMs) were utilized at both sentence and article levels (Qian et al., 2017). Convolutional Neural Networks (CNNs) were applied at various levels, including characters, words, and N-grams (Ruder et al., 2016; Shrestha et al., 2017). Moreover, syntax-augmented CNN models (Zhang et al., 2018), Convolutional Siamese Networks (Saedi and Dras, 2021), and attention-based Siamese Networks (Boenninghoff et al., 2019a) were explored. Additionally, combinations of convolutions and transformers have been employed to learn embeddings for comparison tasks (Andrews and Bishop, 2019).
@@
Barlas and Stamatatos (2020) found that BERT performed well with large vocabularies, outperforming multi-headed RNNs. Fabien et al. (2020) fine-tuned a BERT model for authorship attribution, showing that including additional stylometric and hybrid features in an ensemble model can improve performance. Rivera-Soto et al. (2021) concluded that topic diversity and dataset size are crucial for effective cross-domain transfer. Adaptation through style transfer has not resolved cross-domain issues (Boenninghoff et al., 2019b; Wegmann et al., 2022). Techniques like slanted triangular learning rates and gradual unfreezing can be used to avoid catastrophic forgetting during fine-tuning (Howard and Ruder, 2018).
@@
The incorporation of LLMs in authorship attribution addresses several limitations of traditional methods. Unlike BERT-based models, which require computationally expensive fine-tuning and large amounts of domain-specific data for optimal performance, LLMs can generalize across various domains without fine-tuning, thereby mitigating issues related to domain specificity (Barlas and Stamatatos, 2020). LLMs are also effective with shorter texts, reducing the necessity for long inputs to derive meaningful representations (Eder, 2015). Another key advantage of LLM-based approaches is their ability to provide natural language explanations for their predictions, enhancing transparency compared to hidden text embeddings (Huang et al., 2024). This versatility marks a step forward in overcoming challenges related to data, domain specificity, text length requirements, and explainability faced by earlier methods.
@@ 3. LLM-generated Text Detection
LLMs excel at generating fluent and coherent text, raising concerns about the authenticity and originality of authorship. Detecting LLM-generated text is crucial for several applications, including combating misinformation and disinformation on social media (Gambini et al., 2022; Stiff and Johansson, 2022; Chen and Shu, 2024b), identifying spam (Jindal and Liu, 2008), preventing phishing attacks (Hazell, 2023), identifying fake reviews (Salminen et al., 2022), and detecting machine-generated scientific papers (Rodriguez et al., 2022; Liu et al., 2023a). As a result, the detection of LLM-generated text has garnered significant attention (Kumarage and Liu, 2023; Tang et al., 2023; Wu et al., 2023b; Yang et al., 2023b).
@@ 3.1. Problem Definition
LLM-generated texts are included within the scope of machine-generated texts [^1] Machine-generated texts is also referred to as machine-authored, AI-generated, neural-generated, deepfake text, neural text, or synthetic text.. Machine-generated texts encompass any text produced by automated systems, including simpler language models or rule-based systems (Uchendu et al., 2021). This paper focuses specifically on LLM-generated texts. The task of detecting LLM-generated text involves identifying and distinguishing text created by LLMs from that written by humans. Typically, this task is approached as a binary classification problem (Zellers et al., 2019; Solaiman et al., 2019; Jawahar et al., 2020; Fagni et al., 2021; Mitchell et al., 2023).
@@ 3.2. Methodologies
Evaluating the quality of machine-generated excerpts has traditionally relied on human judgment, which is considered the gold standard for open-domain generation systems (van der Lee et al., 2019; Gehrmann et al., 2019). However, distinguishing between LLM-generated and human-written texts poses significant challenges for humans (Dugan et al., 2023). For example, untrained human reviewers often fail to distinguish GPT-3-generated text from human-written text, identifying it correctly only at a rate consistent with random chance (Clark et al., 2021). Liu et al. (2023a) found that even experienced faculty members and researchers could only achieve about a 50% success rate in identifying GPT-generated academic writings. In contrast, detection algorithms frequently outperform human reviewers in this task (Ippolito et al., 2020).
@@
LLM-generated text detectors can be categorized into metric-based and model-based methods (He et al., 2023), which are further divided into feature-based, neural network-based, zero-shot-based, and watermark-based methods. These detectors are also classified as either white-box or black-box, depending on their access to LLM weights (Tang et al., 2023; Yang et al., 2023b). Watermarking-based methods typically fall under white-box detection, while proprietary models are restricted to black-box methods.
@@
Zero-shot detection methods are generally statistics-based, enabling the identification of LLM-generated text without additional training (Su et al., 2023b). Various statistical measures have been employed for detection, including entropy (Lavergne et al., 2008; Gehrmann et al., 2019), perplexity (Beresneva, 2016; Hans et al., 2024), average log-probability score (Solaiman et al., 2019), fluency (Holtzman et al., 2020), and Zipf’s word frequency law (Zipf, 2016; Piantadosi, 2014) [^2] The frequency of a word decreases as its rank in a frequency-ordered list increases.. Additional methods leverage n-grams (Yang et al., 2023a), Uniform Information Density (UID) (Venkatraman et al., 2023), log rank information (Su et al., 2023b), and various linguistic features such as part-of-speech determiners, conjunctions, auxiliary relations, vocabulary, and emotional tone (Joulin et al., 2017; Tang et al., 2023; Gehrmann et al., 2019).
Various zero-shot detection methods provide distinct strategies to enhance both detection accuracy and efficiency. Gallé et al. (2021) introduced an unsupervised method that identifies over-appearance of repeated higher-order n-grams, distinguishing them from human-generated text. DetectGPT (Mitchell et al., 2023) relies on the observation that LLM-generated passages often fall into regions of negative curvature in log probability. (Bao et al., 2023) enhanced this approach and proposed Fast-DetectGPT, which increases efficiency by using conditional probability curvature. Additionally, some methods leverage LLMs themselves for text classification (Zellers et al., 2019; Solaiman et al., 2019). Different decoding methods are often applied to generate more diverse and less repetitive text, though these can also lead to hallucinations and less verifiable content (Shakeel and Jain, 2021; Guo et al., 2023). Fact-checking methods can mitigate these issues (Zhong et al., 2020; Schuster et al., 2020). Additionally, Krishna et al. (2024) developed a detector that uses information retrieval to store LLM outputs in a database and search for semantically similar content to identify LLM-generated text, though this method raises privacy concerns regarding the storage of user conversations.
@@
Further advancements have improved the robustness, efficiency, and stealthiness of watermarking methods (Hou et al., 2023; Wu et al., 2023a; Zhao et al., 2023). However, there is a trade-off between watermark effectiveness and text quality, as more reliable watermarks require more extensive text modifications (Sadasivan et al., 2023). Additionally, watermarking presents challenges for proprietary LLMs and third-party applications due to the necessity of accessing the language model’s logits (Kirchenbauer et al., 2023). Watermark-based detection methodologies are also vulnerable to paraphrasing attacks (Sadasivan et al., 2023; Krishna et al., 2024).
@@ 3.3. Open Challenges
Detecting LLM-generated texts is challenging due to their high quality and ability to mimic human writing styles. LLMs can emulate human writing so closely that they pose significant challenges to traditional stylometric techniques. They can incorporate complex narrative structures and varied vocabularies, making it difficult to distinguish between human and LLM-generated texts. The rapid evolution of LLMs further complicates detection, as newer versions exhibit different stylistic traits, making detection models quickly obsolete (Chakraborty et al., 2023b). LLM-generated text detectors often struggle to generalize to unseen domains encountered during training (Pu et al., 2023b; Rodriguez et al., 2022; Li et al., 2023a) and tend to perform better on LLMs they were specifically trained on (Pu et al., 2023b; Chakraborty et al., 2023b; Li et al., 2023a).
Existing detectors also lack robustness to various factors, such as alternative decoding strategies (Ippolito et al., 2020), input sequence length (Solaiman et al., 2019), different prompts (Kumarage et al., 2023; Lu et al., 2023), repetition penalties (Fishchuk and Braun, 2023), and human edits (Zhang et al., 2024). Additionally, detectors are vulnerable to adversarial attacks, including homoglyph attacks (Gagiano et al., 2021; Macko et al., 2024), whitespace insertion (Cai and Cui, 2023), syntactic perturbations (Bhat and Parthasarathy, 2020), synonym replacement (Kulkarni et al., 2023), and paraphrasing (Krishna et al., 2024; Shi et al., 2024; Becker et al., 2023).